PITMinutes 07-18-2016

From CASA Wiki
Jump to: navigation, search


R. Bachimanchi, M. Drury, R. Geng, C. Hovater, A. Hutton, G. Krafft, A. Kimber, C. Mounts, R. Nelson, T. Plawski, T. Powers, C. Reece, T. Reilly, R. Rimmer

Agenda and Notes

1. Operations Status and Summary

Microphonic damping feet on 0L04

Correlation studies: systematically torque feet and measure microphonics.

Microphonic stabilization equipment installed on SL-24.

Kirk Davis and Tom taking before and after data.

Idea to modify connections and check mode frequencies. Minimize transfer of microphonics into cavities. Also installed tuner damper to damp MP .

Tuner damper rings in FEL04 also.

Controls improvements to improve recovery times implemented. Better diagnostics Periodic trips being observed again. Snooping around for sources.

RF Software improvement? Capture trips and reduce interactions tuners and MP they generate.

There might be benefit by raising press in SL by 3-4 mbar.

2. C50-12 Warm Girder Study Plan

Both girders out, module next few days.

Sampling plan for girders: about 20 different points.

Samples will be also taken before/after gate valves replaced.

No visible dirt in beam pipe but dust has been observed.

Warm valve problems and failures: due to radiation?

Added valve update and replacement to work.

3. C75 on hold (Gigi Report)

Complete Efforts for 2 prototypes. After report will chart a course forward.

2-3 weeks to complete mockup. 3.5 times 10^7 for Q_L

Several HOM loads located and low loss window located.

Magnetic shielding and hygeine produced good result at cavity locations in CM.

C75 or C50 next year? Waiting test results to decide.

4. Effects of desorbed gas on FE (Geng)

Desorbed gas report coming.

Covering of miniconflat by simple Al foil will not be good enough going forward. Particles can still get around.

Joint goal is to have final written procedure leading to low particulate movements.

NeG pump looking good. Will probably use in warm region.